# PSAppendix - Selected Primary Studies

### v1

@@ 3.2.1. Automatic Search
To identify the primary studies from both academic and gray literature, two different search strings were utilized on the databases, namely IEEEXplore Digital Library, ACM Digital Library, Web of Science, Scopus for academic literature search and Google Search, and Arxiv for gray literature search. Search strings evolve through an exploratory search process. To ensure that key studies are identified, we refined search strings by including relevant keywords (e.g., continuous test, continuous release, and mlsecops). Furthermore, we did not use restricting keywords (e.g., technologies, challenges, and tools) to ensure the relevant studies are captured through the automatic search. Moreover, the automatic search process was performed by two authors to avoid researcher bias. Our search strings are:
@@ 3.2.2. Primary Study Selection from Academic Literature
As we searched the academic literature across multiple databases, some studies were duplicated. At the beginning of the selection process, we removed those duplicated items from our pool of studies, which reduced the number of academic studies from 5519 to 3977. Afterwards, we applied a two-step study selection process to the academic literature: 1) metadata-based selection and 2) full text-based selection. The metadata-based selection is conducted by considering only the paper title, paper keywords, and abstract of the paper, whereas the full-text selection is carried out by reviewing the full text of studies.
During the metadata-based selection, we followed the inclusion and exclusion criteria listed in Table 4. On the other hand, during the full text-based selection, we assessed the quality of each study considering the research aim, research design, methodology, results, limitations of the study, and the value added to research or practice by the study. A more detailed list of quality assessment criteria can be found in our online appendix (onl, 2024). These criteria are adopted from (Barbara Kitchenham, 2007) and (Garousi et al., 2019), discussed and agreed by all authors.
@@
As the final step in academic literature selection, we applied forward and backward snowballing based on the guidelines by Wohlin et al. (Wohlin, 2014). We received 39 more studies as a result of snowballing on the selected academic studies. Following the same inclusion and exclusion, and quality criteria we used in the prior steps, we included 18 out of the 39 studies found by snowballing. Figure 1 depicts a summary of the selection process with the number of papers selected at each step.
Table 3. Inclusion and Exclusion Criteria
Academic literature Gray literature
Inclusion criteria
I1: Full text available and peer-reviewed articles I1: Freely accessible full text articles
I2: Studies that relate to any phase of the ML systems production
Exclusion criteria
E1: Studies that are duplicate
E2: Studies that are written in a language other than English
E3: Studies that are not free to access the full article
E4: Studies that only focus on enhancing machine learning models
E5: Secondary studies (e.g., reviews, surveys, and mapping studies) E5: Videos, audios, question-answering websites, theses
E6: Studies do not present significantly new knowledge or insights
Table 4. Quality Assessment Criteria for Gray Literature
Category Criteria
Authority of the producer Author or publishing authority is reputable or has expertise in the domain
Author or publishing authority has published other relevant articles
Methodology The source has a clearly stated aim
The source is supported by authoritative, contemporary references
Date The item has a clearly stated date
Position related sources Key related gray literature or formal sources have been linked/discussed
Novelty The item enriches or adds something unique to the research
The item strengthens or refutes a current position
Outlet type Gray literature category (i.e., Tier 1, 2 or 3)
Tier 1 - High credibility: Books, magazines, reports, white papers
Tier 2 - Moderate credibility: Annual reports, news articles, presentations, videos, Q/A sites, wiki articles
Tier 3 - Low credibility: blogs, emails, tweets
@@ 3.2.3. Primary Study Selection from Gray Literature
First, we collected a list of studies by applying our search string on Google Search and Arxiv. We stopped our search on Google when the results reached theoretical saturation (Garousi et al., 2019), i.e., meaning that no new content related to MLOps was found. The gathered results roughly covered 50% of all search results. At the end of the study collection, we had a total number of 453 studies including search engine results and pre-prints. Second, we applied the inclusion and exclusion criteria reported in Table 4 to select primary studies for our work. We removed video contents, contents that require non-free access, and duplicate studies. We ended up having 423 studies. Third, we applied our quality assessment criteria, listed in Table 4 which are adopted from (Garousi et al., 2019), on the remaining studies. We iteratively simplified the quality criteria provided by Garousi et al. (Garousi et al., 2019) to ensure we captured all the relevant content. For example, we excluded many of the methodology related criteria as MLOps is still an emerging field and many are still trying to define it and its associated technologies and concepts. Hence studies tend to skip reporting experimental data or methodology. After the quality assessment, we selected studies that met at least seven out of nine criteria listed in the table.
Similar to the academic study selection process, we established consensus for gray literature selection by having two authors independently review the first 50 of the search engine results, and comparing their results with each other at the beginning of the quality assessment. Later, two authors separately continued to review the remaining studies and ended up with 41 primary studies. As the last step, a snowballing process was performed on the selected gray literature items. 27 more studies were collected from snowballing and 19 of them were selected as primary studies.
@@ 3.3. Data Extraction and Thematic Analysis
Thematic analysis is a qualitative data analysis method that is widely used in systematic literature reviews (Dissanayake et al., 2022; Croft et al., 2022). For synthesizing and analyzing the qualitative data extracted from the reviewed studies, we followed the six-phase process of the thematic analysis as per the guidelines provided by Braun et al. (Braun and Clarke, 2006).
As the first step of the process, selected articles were divided among three authors and each author familiarised them-self with data by reading the full-text and identifying content related to our research questions using a predefined data-extraction form (Garousi et al., 2019; Barbara Kitchenham, 2007). To ensure the inter-rater agreement regarding the data extraction process, a pilot data extraction was carried out on six papers by each author and results were discussed. Later, data extraction was carried out separately by splitting the primary studies among all the authors. As the second step, initial codes were generated ensuring that all data extracts from the primary studies are coded and collated together within each code. We used a shared spreadsheet as a Code Book to record codes and their definitions. This allowed authors to coordinate, validate the generated codes, and re-use codes when applicable. We also maintained an Obsidian  [^1] https://obsidian.md/ vault containing a collection of notes, where each note was created per generated code along with related data extracts from the primary studies. These notes collated data extracts to generated codes. In the third step, we analyzed the generated codes and combined them to generate a collection of candidate hierarchical themes (themes, sub-themes, and sub-sub-themes). Three authors carried out this step, where each author generated the themes related to one of the research questions. The generated themes were reviewed by the other two authors to refine and validate them. The fourth step is for defining the final names and definitions of the refined themes followed by the final step of reporting the synthesized data.
3.4. Overview of the Selected Studies
There are a total number of 220 primary studies selected from both the academic and gray literature; 150 of them are selected from the academic literature while 48 of them correspond to the gray literature. Figure 2 reports the distribution of studies over time and according to their literature types. The publication year distribution of our selected studies shows that MLOps-focused studies have been increasing each year since 2019. Note that, the year 2023 column shows the statistics of studies published up to September 2023.
Figure 2. Publication counts by year and literature type
4. RQ1: How do researchers and practitioners perceive MLOps?
@@
MLOps is often defined by the development, deployment, and operationalization of ML systems, as well as the application of cultural practices in ML projects to enable productionalization of ML systems. The most common definitions that occur in the literature are continuous ML, DevOps for ML, and automation of ML pipelines. In the following subsections, we define MLOps in more detail, from the most common definitions to less common ones, e.g., 34 studies define MLOps as “continuous ML”, 32 studies define it as “DevOps for ML”, 18 studies refer to it as cultural and organizational practices for ML. We provide the full statistics in our online appendix (onl, 2024).
4.1.1. Continuous ML
@@
Continuous ML draws from the well-known software engineering practice CI/CD (continuous integration and delivery) which aims to integrate the software frequently to deliver high-quality products (see Appendix B.1). This concept takes place in building ML systems as well, and it is often defined by the combination of CI/CD with continuous training (CT) [80]. MLOps brings continuous practice into many phases of ML life-cycle [95] (i.e., model design [5], model training [1], serving [67], validating [86], monitoring [5], model delivery, [28] and integration of pipelines [102]). MLOps aims at rapid development and operations through continuous practices, adapting existing solutions quickly to changing conditions (e.g. business needs, concept shift) [87], [85], [102]. Moreover, continuous practices allow feedback loops between different phases that are critical in ML projects due to the experimental nature of ML engineering. For example, ML models are often updated after the initial phase of monitoring, and feature engineering and model serving tasks are repeated with different model needs [10].
4.1.2. DevOps for ML
@@
MLOps originates from the DevOps concept [1], [29], [32],[141] and targets development (ml) and operationalization (ops) of ML systems [30], [88], [102]. MLOps adopts practices and culture associated with DevOps [32], [85], [156] aiming to standardize streamline development, deployment, operation, and management of ML systems [108], [115]. MLOps unifies development and operational cycles carried out by different roles, to ensure automated, frequent, and continuous delivery of high-performing ML systems [67].
4.1.3. Automation of ML pipelines
@@
Another highlighted aspect of MLOps in the literature is the automation of pipelines. Automation leads to more rapid development and deployment [28], and easier management of the infrastructure that ML models are served [72]. MLOps advocates automation [102] at all steps of ML system development and operations including activities from data collection, model building and deployment, and monitoring [1], [17], [30]. Automation eases the deployment by providing a more simple release and environment configuration process, thus, researchers can focus more on model optimization [44].
4.1.4. Cultural and organizational practices to deliver ML systems
@@
Developing and operationalizing an ML-enabled system necessitate collaboration between many different roles, e.g., business experts, ML engineers, and software engineers. As the number of ML applications and their complexity grows in the community, the need for establishing a standard for new roles and responsibilities, and organizational collaboration practices becomes more prevalent. MLOps is reported as a set of tools and organizational practices that lead to more sustainable development and operationalization of ML-enabled systems [70], through enabling collaboration across diverse teams, e.g., data science team and IT professionals [45].
4.1.5. Intersection of Data Science and ML
@@
As emphasized by several studies [81], [87], [86], [96], MLOps shares a similar philosophy with DevOps, but has unique aspects related to data and model. Hence, definitions of MLOps point out that MLOps is the intersection of data engineering, data science, machine learning, DevOps, and software engineering [157]. Further, MLOps is closely linked with agility in data science [30], [32], [36], [83]. Specifically, producing and releasing ML models in an iterative manner is the application of agile principles (Fowler et al., 2001) in ML projects.
4.1.6. ML system productionalization
One prominent definition of MLOps is the process of putting ML solutions in production [7]. Productionalization of ML refers to serving the ML solution to customers in a live environment, which involves tasks and challenges different than development. MLOps targets minimizing the leap between experimental development and production environment [163], and transferring the developed ML model to production at scale by ensuring reproducibility, reliability, and efficiency [71], [23]. Thus, MLOps involves practices to unify the development, deployment, and operationalization of ML systems through processes, tools, automation, and repetition to create a streamlined workflow from development through operationalization [22], [41], [89], [96].
4.1.7. ML life-cycle management
Another commonly emphasized definition of MLOps is the management of life-cycle, which refers to the end-to-end processes of development, deployment, and monitoring [37], [44], [54], [61] These life-cycle phases work in flows and interact with each other through triggers, loops, and common artifacts [61]. Further, serving ML solutions requires integration with middleware, software systems and services [2], ensuring fairness of predictions [152], scaling ML solution accordingly to infrastructure (e.g., decentralized development [64], MLOps on the edge [144]). In addition to that, the nature of ML models includes many complexities, some of which are noise in data, model tuning, and large amounts of artifacts. MLOps is the management, standardization and optimization of all these complexities considering reproducibility, stability, optimization, and scalability of pipelines [91], [94], [95], and managing various artifacts generated over the life-cycle of a project (i.e., multiple versions of experiments, datasets, models, and metadata related to models) [144].
4.1.8. Unify of ML development, deployment, and operations
@@
Figure 3 shows a comprehensive MLOps pipeline that includes MLOps activities. We elaborate more on these activities in the following subsections.
Figure 3. A pipeline for ML development and operations
@@ 4.2.1. ML project initiation
ML project initiation mainly includes activities such as business analysis, domain understanding, planning the project, specifying system requirements, and designing the ML system. Many stakeholders take part in these activities, i.e., business analysts, business owners, domain experts, data scientists, and ML and MLOps engineers.
Specification of business and system requirements is crucial for the planning and execution of an ML project, and for the successful deployment and operation of the ML system as well [105], [111]. One of the initial steps in identifying business requirements is the formulation of the business problem to establish a clear idea of the problem to be addressed with the ML system [111], [41]. Formulation of the business problem helps the team evaluate the business value to be gained by solving the problem [111], analyzing whether the problem could be addressed by building an ML system, and the feasibility of engineering of an ML system [116]. Identification of business requirements includes further steps such as obtaining security objectives of the company, e.g. confidentiality and privacy requirements of the data used through ML system [105], and regulatory rules [99].
Planning an ML system also includes tasks related to software project management knowledge areas (Bourque and Fairley, 2004); i.e., determining the scope of the project [90], specifying ML use cases [70], designing the model architecture, pipeline, and infrastructure of the system [90], [70], identification of performance metrics that will be used to assess the ML model [111], determining how statistics will be collected during the monitoring phase and how the feedback of business owners will be received [70].
4.2.2. Data collection
Data collection is the first step in the ML pipeline, referring to the acquisition of raw data that will be processed in later steps to train a ML model. Often, raw data is gathered from relevant data sources, e.g., sensors [24], GitHub code repositories [27], or cameras [34].

### v2

@@ 2.2.1. Automatic Search
For academic literature, we searched IEEEXplore Digital Library, ACM Digital Library, Web of Science, and Scopus using complex Boolean queries incorporating all related keywords into the search string to maximize the retrieval of relevant content from these databases. For gray literature, we used simpler, flexible search string as complex queries would limit the results of the search engine.
We developed and evolved search strings through an exploratory search process, incorporating relevant keywords (e.g., continuous test, mlsecops) while avoiding restricting keywords (e.g., technologies, challenges, and tools) to capture relevant content. Based on this process, we reached two different search strings for peer-reviewed and gray literature (Table 2).
Table 2. Search Strings
@@
The search process was independently executed by two authors. Individual results were combined in a shared file, and non-overlapping items—potentially due to location-based differences—were reviewed and discussed for inclusion.
@@ 2.2.2. Primary Study Selection from Academic Literature
We removed those duplicated items from our pool of studies, which reduced the number of academic studies from 5519 to 3977. Afterwards, we applied a two-step study selection process to the academic literature: (1) metadata-based selection and (2) full text-based selection. The metadata-based selection is conducted by considering only the paper title, paper keywords, and abstract of the paper, whereas the full-text selection is carried out by reviewing the full text of studies.
During the metadata-based selection, we followed the inclusion and exclusion criteria listed in our online appendix (onl, 2024). On the other hand, during the full text-based selection, we assessed the quality of each study considering the research aim, research design, methodology, results, limitations of the study, and the value added to research or practice by the study. A more detailed list of quality assessment criteria can be found in our online appendix (onl, 2024). These criteria are adopted from (Barbara Kitchenham, 2007) and (Garousi et al., 2019), discussed and agreed by all authors.
@@
As the final step in academic literature selection, we applied forward and backward snowballing based on the guidelines by Wohlin et al. (Wohlin, 2014). We received 39 more studies as a result of snowballing on the selected academic studies. Following the same inclusion and exclusion, and quality criteria we used in the prior steps, we included 18 out of the 39 studies found by snowballing, creating a final set of 150 peer-reviewed literature.
@@ 2.2.3. Primary Study Selection from Gray Literature
First, we collected a list of studies by applying our search string on Google Search and Arxiv. We stopped our search on Google when the results reached theoretical saturation (Garousi et al., 2019), i.e., meaning that no new content related to MLOps was found. The gathered results roughly covered 50% of all search results. At the end of the study collection, we had a total number of 453 studies including search engine results and pre-prints. Second, we applied the inclusion and exclusion criteria (onl, 2024) to select primary studies. This involved removing video contents, non-free materials, and duplicates already identified through academic search. We ended up having 423 studies. Third, we applied our quality assessment criteria, reported in our online appendix (onl, 2024), which are adopted from (Garousi et al., 2019), on the remaining studies. We iteratively simplified the quality criteria provided by Garousi et al. (Garousi et al., 2019) to ensure we captured all the relevant content. For example, we excluded many of the methodology related criteria as MLOps is still an emerging field and many are still trying to define it and its associated technologies and concepts. Hence studies tend to skip reporting experimental data or methodology. After the quality assessment, we selected studies that met at least seven out of nine criteria listed in the table. As the last step, a snowballing process was performed, leading to a final set of 48 gray literature.
@@ 2.3. Data Extraction and Thematic Analysis
Thematic analysis is a qualitative research methodology widely used in systematic literature reviews to understand patterns within the extracted data by identifying, analyzing, and reporting themes (Dissanayake et al., 2022; Croft et al., 2022). The thematic analysis begins with researchers familiarizing themselves with the data to gain a comprehensive understanding. Based on this understanding, they systematically create and assign codes to the data to signify interesting features and potential themes. The codes are then sorted into themes, which capture significant patterns in the data. Finally, themes are reviewed, accompanied by clear definitions and names for reporting. We followed the guidelines for thematic analysis established by Braun et al. (Braun and Clarke, 2006) to carry out the analysis.
Understanding: The process commenced with all three authors engaging in data extraction to facilitate an in-depth understanding of the source material. We performed a pilot data extraction on six papers. To enhance inter-rater agreement, we discussed the results to refine the data extraction process and the predefined data-extraction form (Garousi et al., 2019; Barbara Kitchenham, 2007). Subsequently, data extraction was carried out separately, with the primary studies divided among the authors.
Coding: The coding process began after the completion of data extraction. We utilized a shared spreadsheet to maintain an up-to-date Code Book. Additionally, we maintained an Obsidian  [^1] https://obsidian.md/ vault to supplement the details in the Code Book within the spreadsheet. This vault contains a collection of notes, each providing a detailed description of a code along with related data extracted from the primary studies. These artefacts were shared and version-controlled, enabling authors to coordinate, validate the generated codes, and reuse codes when applicable. Each of the three authors focused on coding for one specific research question, with biweekly catch-up meetings to synchronize our efforts.
Theme generation: We analyzed the generated codes and combined them to create a hierarchy of candidate themes (including themes, sub-themes, and sub-sub-themes). Each author independently generated themes corresponding to their coded research question. The other two authors then reviewed the generated themes for validation and refinement. Finally, we defined the final names and definitions of the refined themes before reporting the results.
2.4. Threats to Validity
Potential threats to the validity of our study primarily concern the processes of study search, study selection, data extraction, and theme analysis. We collected studies from well-known academic databases and general search engines to ensure a comprehensive selection of the studies. Overview of selected studies and their distributions over time is available at our online appendix (onl, 2024). While going through search results, we excluded non-free resources, which is a limitation of our study. Additionally, we applied backward and forward snowballing techniques to identify potentially missed studies. We designed search strings to ensure broad coverage while avoiding restrictive keywords. Complex queries including many keywords were used in academic databases, while simpler queries were applied for gray literature to accommodate search engine limitations. This approach minimized the risk of omitting critical studies while maintaining relevance to MLOps. To ensure consistency in applying inclusion/exclusion criteria and quality assessment, we conducted pilot evaluations. All authors independently assessed randomly selected studies and discussed results to align on terminology and evaluation standards. In cases of uncertainty, decisions were made through consensus, ensuring reliability in the quality assessment process. We used a standardized template to extract key study details (e.g., objectives, methodologies, findings) (onl, 2024). A pilot extraction phase helped align understanding, with discrepancies resolved through discussion. For theme analysis, we employed an iterative coding process where three authors independently coded the data, resolving disagreements through discussion.
3. RQ1: Conceptualization of MLOps
@@
Eight types of definitions for MLOps have been identified in the existing literature. These definitions can be considered the conceptual foundation for MLOps work, as they encompass what authors of the work perceive to be the primary challenges in ML productionalization, thereby delineating the scope of MLOps solutions (e.g., best practices, tools) to be proposed.
The following subsections introduce the identified types of MLOps definitions, ranging from the most common to the least common. For example, 34 studies define MLOps as “continuous ML,” 32 refer to it as “DevOps for ML,” and 18 studies describe it as encompassing cultural and organizational practices for ML. We provide the complete statistics in our online appendix (onl, 2024). Figure 2 visualizes the relationship between eight definition types based on the types of ML productionalization problems on which they focus.
Figure 2. MLOps definitions and the ML productionalization problems they address
3.1.1. MLOps-as-Continuous-ML
@@
This definition considers MLOps as the integration of continuous engineering practices (e.g., Continuous Integration (CI) and Continuous Delivery (CD)) into the lifecycle of ML-enabled software systems. The goal of MLOps-as-Continuous-ML is to enable rapid development and operations through continuous practices, allowing for the swift adaptation of existing solutions to changing conditions (e.g., business needs, concept shifts) [87], [85], [102]. Moreover, continuous practices facilitate feedback loops between different phases critical in ML projects due to the experimental nature of ML engineering. For example, ML models are often updated after the initial monitoring phase, and feature engineering alongside model serving tasks are repeated to accommodate varying model needs [10].
The concept of MLOps-as-Continuous-ML is inspired by the well-known software engineering practice of CI/CD (continuous integration and delivery), which aims to integrate software to deliver high-quality products frequently (see Appendix B.1), particularly emphasizing the parallels between delivering software to production and delivering ML to production. Continuous training (CT) [80] is often introduced as the ML-specific element. MLOps-as-Continuous-ML integrates continuous practices into many phases of the ML lifecycle [95], including model design [5], model training [1], serving [67], validating [86], monitoring [5], model delivery [28], and the integration of pipelines [102].
3.1.2. DevOps-for-ML
This definition considers MLOps as an extension of DevOps practices for ML. It should be noted that while continuous engineering practices are a key and perhaps the most well-known element of DevOps, the scope of DevOps itself extends beyond this to encompass practices such as streamlining infrastructure provisioning via Infrastructure-as-Code, facilitating data-driven decision-making by providing developers with on-demand observability, as well as fostering the cultural shifts necessary to enable these cross-silo activities. These cultural shifts might include promoting collaboration between development and operations teams, encouraging a mindset of continuous improvement, and embracing automation.
@@ 3.1.2. DevOps-for-ML
MLOps originates from the DevOps concept [1], [29], [32], [141] and targets the development (dev) and operationalization (ops) of ML systems [30], [88], [102]. MLOps adopt practices and cultural aspects associated with DevOps [32], [85], [156], aiming to standardize and streamline the development, deployment, operation, and management of ML systems [108], [115]. MLOps unifies the development and operational cycles carried out by different roles, ensuring the automated, frequent, and continuous delivery of high-performing ML systems [67].
3.1.3. MLOps-as-Pipeline-Automation
@@
This definition considers MLOps as the practice of automating the repetitive activities in the ML lifecycle by constructing and managing automated pipelines. Details regarding ML lifecycle activites would be discussed in section 3.2. Some benefits of automation include more rapid development and deployment [28] and easier management of the infrastructure to which ML models are deployed [72]. MLOps advocates for automation [102] at all stages of ML system development and operations, including activities such as data collection, model building and deployment, and monitoring [1], [17], [30]. Automation simplifies deployment by providing a more straightforward release and environment configuration process, allowing researchers to focus more on model optimization [44].
3.1.4. MLOps-as-Culture
@@
Developing and operationalising an ML-enabled system necessitates collaboration between many different roles, e.g., business analysts, ML engineers, and software engineers. As the number of ML applications and their complexity grows in the community, the need for establishing a standard for new roles and responsibilities and organisational collaboration practices becomes more prevalent. This definition considers MLOps as the set of tools and managerial practices that lead to more sustainable development and operationalisation of ML-enabled systems [70] through enabling collaboration across diverse teams, e.g., data science team and IT professionals [45].
3.1.5. MLOps-as-Agile-Data-Science
@@
As emphasized by several studies [81], [87], [86], [96], MLOps shares a similar philosophy with DevOps, but has unique aspects related to data and model. Further, MLOps is closely linked with agility in data science [30], [32], [36], [83]. Specifically, producing and releasing ML models iteratively is the application of agile principles (Fowler et al., 2001) in ML projects. Hence, this definition describes MLOps as the intersection between data engineering, data science, machine learning, and operation to facilitate an agile end-to-end practice.
3.1.6. MLOps-as-ML-Productionalisation
This definition considers MLOps as the process of operating ML models, which involves bringing them into a production environment [7]. The purpose of MLOps from this perspective is to minimize the gap between experimental development and the production environment [163], as well as to facilitate the transfer of developed ML models to production at scale by ensuring reproducibility, reliability, and efficiency [71], [23]. Research based on this definition of MLOps focuses on the processes, standards, and tools for packaging and integrating ML models into software systems, as well as on streamlined workflows from development to operationalization [22], [41], [89], [96].
3.1.7. MLOps-as-ML-Life-Cycle-Management
This definition considers MLOps as the set of practices and tools to manage lifecycle activities of MLOps and their corresponding artefacts to ensure reproducibility, stability, optimization, and scalability. One focus of MLOps research from this perspective is the handling of triggers, loops, and common artefacts that link various MLOps lifecycle activities [61], such as managing multiple versions of experiments, datasets, models, and metadata related to models [144]. Another focus is the management, standardization and optimization of workflows across lifecycle activities [91], [94], [95]. Another focus is permeating the lifecycle with responsible practices such as fairness of predictions [152].
3.1.8. MLOps-as-a-Unified-Engineering-Practice
@@
Figure 3 shows a comprehensive MLOps pipeline that includes core MLOps activities and roles associated with them. We elaborate more on the activities in this section and roles in section 3.3.
Figure 3. A pipeline for ML development and operations, and associated roles: Domain Experts (DME); Business Analysts (BA); Data Engineers (DE); Data Scientists (DS); ML Engineers (MLE); MLOps Engineers (MLOE); DevOps Engineers (DOE); Software Engineers (SE); Quality Assurance Engineers (QAE); Operational Engineers (OPE); Technical and Solution Architects (ARC); Security Experts (SE); Business Owners/Managers (BM); Product Owners/Managers/Experts (PM); Data Governance Officer (DGO), Data Quality Board (DQB), Data Steward (DST)
@@ 3.2.1. ML project initiation
ML project initiation mainly includes activities such as business analysis, domain understanding, project planning, specifying system requirements, and ML system design. Many stakeholders take part in these activities, i.e., business analysts, domain experts, data engineers and scientists, and ML and MLOps engineers. Specification of business and system requirements is crucial for planning and execution of an ML project, and successful deployment and operation of the ML system [105], [111]. The process begins with formulating the business problem to establish a clear objective [41], helping assess business value [111], determine the feasibility of engineering an ML system [116]. Further steps include obtaining security objectives of the company, e.g., confidentiality and privacy requirements of the data used through ML system [105], and regulatory rules [99]. Moreover, ML system planning incorporates software project management knowledge areas (Washizaki, 2024) such as defining project scope [90], specifying ML use cases [70], designing model architecture, pipeline, and infrastructure [90], [70], identifying performance metrics for model evaluation [111], and establishing monitoring strategies with feedback mechanisms for business owners [70].
3.2.2. ML lifecycle activities